OpenStudy (anonymous):

Find the area of triangle ABC if a=12 b=14 c=17

OpenStudy (nikita2):

there is Geron's formula...

OpenStudy (nikita2):

Actually Heron's formula; S = sqrt(p(p-a)(p-b)(p-c)), where p = (a+b+c)/2.

OpenStudy (radar):

1. calculate the perimeter P (sum of the three sides). 2. divide that by 2 and call that s. 3. Then use the Heron's formula below: \[A=\sqrt{s(s-a_(s-b)(s-c)}\] where a, b, and c are the length of the triangle legs.

OpenStudy (radar):

\[correction A=\sqrt{s(s-a)(s-b)(s-c)}\]

OpenStudy (radar):

Try it yoltsbrand it should work.